OpenStudy (sangelc):

If 2(x – 5) = –11, then x = ?

OpenStudy (kenrick143):

simpilfy...

thomaster (thomaster):

2x - 10 = -11 +10 +10 2x = -1 /2 /2 x = ??

OpenStudy (ben00):

it would equal a minus half.